OpenStudy (anonymous):

The area of the sector of a circle having a central angle of 120 is 24sq.cm What is the radius of the circle? What is the perimeter of the arc of the sector?

OpenStudy (anonymous):

how many radians is 120 degrees?

OpenStudy (anonymous):

\(A_{sector}=\pi\theta\)

OpenStudy (anonymous):

where \(\theta\) is in radians

OpenStudy (anonymous):

Area of the sector: 1/2pi r^2 [\theta \]

OpenStudy (anonymous):

oh yeah, sorry. length of a segment is what i wrote

OpenStudy (nincompoop):

you already have the area, plug and chug

OpenStudy (anonymous):

wrong reply, sorry :) 24 cm^2 = 1/2pi r^2 120 deg (pi/180)

OpenStudy (anonymous):

\(A_{sector}=\frac{\theta}{2}r^2\)

OpenStudy (anonymous):

then, square root :) you have an answer for no. 1.

OpenStudy (nincompoop):

you're looking for r so isolate it and then square root

OpenStudy (anonymous):

Perimeter of arc sector: r(2+theta)

OpenStudy (nincompoop):

I don't think she's following... LUL it's finals time. we shouldn't even be helping at this point it beats the purpose of learning and it is not fair to those who study arduously and those who fail